For the polynomial below, -2 is a zero.
g(x)=x^3 -2x^2 -11x -6

Express g(x) as a product of linear factors.
g(x)=

Answer by stanbon(75887)

For the polynomial below, -2 is a zero.
g(x)=x^3 -2x^2 -11x -6
---
-2)....1....-2....-11....-6
........1.....-4.....-3...|..0
Quotient: x^2-4x-3
Solve x^2-4x-3 = 0
x = [4+-sqrt(16-4*-3)]/2 = [4+-sqrt(28)]/2
------
x = [4 +- 2sqrt(7)]/2
----
x = 2+sqrt(7) or x = 2-sqrt(7)
==================================
Express g(x) as a product of linear factors.
g(x) = (x+2)(x-2-sqrt(7))(x-2+sqrt(7))
========================================
Cheers,
Stan H.